OpenStudy (anonymous):

x and y intercepts for equation 3y=9x+15

OpenStudy (anonymous):

For finding x intercept, put y = 0 And finding y intercept put x = 0 Can you do it now?/ Firstly put y = 0 to find x..

OpenStudy (anonymous):

I am expecting a reply from you @Rosie123

OpenStudy (anonymous):

What will you get when you will put y = 0 in 3y = 9x + 15..

OpenStudy (anonymous):

can I try @waterineyes

OpenStudy (anonymous):

Thisis one of my lessons too this week. @waterineyes

OpenStudy (anonymous):

Yes you can but do not provide the whole solution.. Suppose you are a teacher and he or she is a student.. Just guide here...

OpenStudy (anonymous):

Go ahead now.

OpenStudy (anonymous):

y = 0 in 3y = 9x + 15 3(0)=9x+15 0=9x+15

OpenStudy (anonymous):

@waterineyes i am here

OpenStudy (anonymous):

15-0=9x+15-15

OpenStudy (anonymous):

Can you solve what nia is saying to you ?? @Rosie123

OpenStudy (anonymous):

i am confussed because i got x intercept is -5/3,0 and y intercept is 0,5

OpenStudy (anonymous):

@waterineyes am I in a right way?

OpenStudy (anonymous):

So you are absolutely right @Rosie123

OpenStudy (anonymous):

oh so if u have to graph it how can u graph -5/3?

OpenStudy (anonymous):

There is one mistake nia.. 9x + 15 = 0 9x + 15 - 15 = 0 -15 15 will go there as negative @nia_maepenix

OpenStudy (anonymous):

thanks a lot @waterineyes

OpenStudy (anonymous):

Practice more of this algebra nia..

OpenStudy (anonymous):

oh so if u have to graph it how can u graph -5/3?

OpenStudy (anonymous):

@waterineyes Thanks Dear!

OpenStudy (anonymous):

What is -5/3 ?? = -1.66 So take two points at (-1.66, 0) and (0,5) and join these two points by a line @Rosie123

OpenStudy (anonymous):

ok i got it thanks a lot u guys :)

OpenStudy (anonymous):

So try to draw a graph now..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@cwrw238 cute drawing.

OpenStudy (anonymous):

i know its cute :)

OpenStudy (cwrw238):

hhmmm ty lol

OpenStudy (anonymous):

yw :)